OpenStudy (anonymous):

A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax and x represents the original price of the item. : C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale? Remember, do not include the $ symbol.

OpenStudy (anonymous):

@HectorH

OpenStudy (anonymous):

@Teddyiswatshecallsme

OpenStudy (david27):

C = 0.765(100) + 0.06[0.765(100)] C = 76.5 + 0.06(76.5) C = 76.5 + 4.59 C = 81.09 After the discount, Chris will pay $81.09 for the coat.

OpenStudy (anonymous):

thnx :)

OpenStudy (david27):

your welcome